L{ e2t (t-1)2 }


1
Expert's answer
2021-09-06T19:15:47-0400

"L[e^{2t}(t-1)^2]=L[e^{2t}(t^2-2t+1)]=\frac{(s-2)^2-2(s-2)+2}{(s-2)^3}=\frac{s^2-6s+10}{(s-2)^3}."